Quantrex Quantrex AcademyJEE · NEET · NDA PYQs with solutions Open app
JEE Advanced 2023Paper 2
Exam
JEE Advanced
Year
2023
Paper
Paper 2
Subject
Mathematics
Chapter
Probability

JEE Advanced 2023 · Paper 2 · Mathematics · Probability

Consider an experiment of tossing a coin repeatedly until the...

Consider an experiment of tossing a coin repeatedly until the outcomes of two consecutive tosses are same. If the probability of a random toss resulting in head is 13\dfrac{1}{3}, then the probability that the experiment stops with head is

Options

  1. A13\dfrac{1}{3}
  2. B521\dfrac{5}{21}
  3. C421\dfrac{4}{21}
  4. D27\dfrac{2}{7}

Correct answer

B. 521\dfrac{5}{21}

Step-by-step solution

Given,

P(H)=13P ( H ) = \dfrac{1}{3}

So, P(T)=1−13=23P ( T ) = 1 - \dfrac{1}{3} = \dfrac{2}{3}

Now probability of the event is given by,

P(E)=P(HH)+P(THH)+P(HTHH)+P(THTHH)+P(HTHTHH)+P(THTHTHH)+…P ( E ) = P ( HH ) + P ( THH ) + P ( HTHH ) + P ( THTHH ) + P ( HTHTHH ) + P ( THTHTHH ) + \ldots

⇒P(E)=13×13+23×13×13+13×23×13×13+435+436+837+838+…⇒ P ( E ) = \dfrac{1}{3} \times \dfrac{1}{3} + \dfrac{2}{3} \times \dfrac{1}{3} \times \dfrac{1}{3} + \dfrac{1}{3} \times \dfrac{2}{3} \times \dfrac{1}{3} \times \dfrac{1}{3} + \dfrac{4}{{3}^{5}} + \dfrac{4}{{3}^{6}} + \dfrac{8}{{3}^{7}} + \dfrac{8}{{3}^{8}} + \ldots

⇒P(E)=132+233+234+435+436+837+838+…⇒ P ( E ) = \dfrac{1}{{3}^{2}} + \dfrac{2}{{3}^{3}} + \dfrac{2}{{3}^{4}} + \dfrac{4}{{3}^{5}} + \dfrac{4}{{3}^{6}} + \dfrac{8}{{3}^{7}} + \dfrac{8}{{3}^{8}} + \ldots

⇒P(E)=(132+234+436+….)+(233+435+837+….)⇒ P ( E ) = \left(\dfrac{1}{{3}^{2}} + \dfrac{2}{{3}^{4}} + \dfrac{4}{{3}^{6}} + \ldots .\right) + \left(\dfrac{2}{{3}^{3}} + \dfrac{4}{{3}^{5}} + \dfrac{8}{{3}^{7}} + \ldots .\right)

⇒P(E)=(132(1−232))+(233(1−232))⇒ P ( E ) = \left(\dfrac{\dfrac{1}{{3}^{2}}}{\left(1 - \dfrac{2}{{3}^{2}}\right)}\right) + \left(\dfrac{\dfrac{2}{{3}^{3}}}{\left(1 - \dfrac{2}{{3}^{2}}\right)}\right)

⇒P(E)=17+221=521⇒ P ( E ) = \dfrac{1}{7} + \dfrac{2}{21} = \dfrac{5}{21}

Practice Probability on Quantrex Academy →

More from Probability

Suppose that Box I contains 6 red balls and 9 green balls, and Box II contains 8 red… JEE Advanced 2026A candidate has to go to the examination centre to appear in an examination. The… JEE Main 2026A bag contains 6 blue and 6 green balls. Pairs of balls are drawn without replacement… JEE Main 2026A bag contains (N+1) coins - N fair coins, and one coin with 'Head' on both sides. A coin… JEE Main 2026The probabilities that players A and B of a team are selected for the captaincy for a… JEE Main 2026A letter is known to have arrived by post either from KANPUR or from ANANTPUR. On the… JEE Main 2026From a month of 31 days, 3 different dates are selected at random. If the probability… JEE Main 2026A coin is tossed 8 times. If the probability that exactly 4 heads appear in the first six… JEE Main 2026A man throws a fair coin repeatedly. He gets 10 points for each head he throws and 5… JEE Main 2026The probability distribution of a random variable X is given below : (array… JEE Main 2026A bag contains 10 balls out of which k are red and ( 10-k ) are black, where 0 ≤ k ≤ 10 .… JEE Main 2026Let S be a set of 5 elements and P(S) denote the power set of S. Let E be an event of… JEE Main 2026From a lot containing 10 defective and 90 non-defective bulbs, 8 bulbs are selected one… JEE Main 2026Bag A contains 9 white and 8 black balls, while bag B contains 6 white and 4 black balls.… JEE Main 2026From the first 100 natural numbers, two numbers first a and then b are selected randomly… JEE Main 2026If a random variable x has the probability distribution array |c|c|c|c|c|c|c|c|c| x & 0 &… JEE Main 2026 Full Probability list Probability JEE Maths PYQsJEE Mathematics PYQs All JEE Advanced PYQs